# Axionlike particle-assisted supercooling chiral phase transition in QCD: Identifying Coleman-Weinberg type-chiral phase transition in QCD-like scenarios

### v2

@@ Abstract
We propose a new scenario to realize the Coleman-Weinberg (CW) type chiral phase transition in the QCD thermal history. This scenario predicts an axionlike particle (ALP) with mass $\gtrsim$ sub GeV, consistently with the current experimental and cosmological bounds. The chiral phase transition is evaluated by monitoring ordinary QCD setup in a view of a two-flavor Nambu-Jona-Lasinio model at the mean-field approximation level. The present work thus can open a new window to search for the ALP associated with the QCD phase transition epoch of the thermal history, to be probed by gravitational wave and primordial black hole productions. This proposal is based on a generic classification of the order of the chiral phase transition in a view of the scale violation classes: the soft-scale breaking term and the CW-type scale anomaly term, in or off medium with or without chemical potentials. On this theoretical ground, we also revisit existing scenarios which undergo the supercooling chiral phase transition, such as nearly scale-invariant QCD and QCD with a large baryon chemical potential.
@@ I Introduction
First order phase transitions or supercooling phase transitions currently involve broader interests over theoretical particle physics and cosmology Mazumdar:2018dfl . A particular motivation can be seen in cosmological applications to the thermal history of the Universe: the first-order/supercooling phase transition serves as productions for stochastic gravitational wave (GW) background and primordial black hole (PBH) formation Jung:2021mku ; Kodama:1982sf ; Hall:1989hr ; Kusenko:2020pcg ; Hashino:2021qoq ; Hashino:2022tcs ; He:2022amv ; Liu:2021svg ; Kawana:2022olo ; Lewicki:2023ioy ; Gouttenoire:2023naa ; Gouttenoire:2023pxh ; Salvio:2023blb ; Conaci:2024tlc ; Banerjee:2023brn ; Banerjee:2023vst ; Banerjee:2023qya ; Lewicki:2024ghw ; Flores:2024lng ; Cai:2024nln ; Florentino:2024kkf ; Goncalves:2024vkj ; Banerjee:2024fam ; Shao:2024dxt ; Banerjee:2024cwv ; Wu:2024lrp ; Hashino:2025fse ; Murai:2025hse ; Franciolini:2025ztf ; Kierkla:2025vwp ; Kanemura:2024pae as nonparticle dark matter candidate Carr:2020gox ; Baldes:2023rqv ; Balaji:2024rvo ; Jinno:2023vnr ; Zhang:2025kbu ; Kanemura:2024pae . It has also been argued that the supercooling electroweak or QCD phase transition can provide a possible initial condition for slow-roll inflation Zhang:2023acu ; Liu:2024xrh , and subsequent particle production Wang:2022ygk ; Wang:2024tda .
@@
One typical modeling to access those cosmological applications would be a scale-symmetry phase transition via the Coleman-Weinberg (CW) mechanism Coleman:1973jx , assuming an appropriate presence of scalar field (presumably allowing coupled to the Standard Model (SM) plasma). All dimensionful parameters are absent at tree level, but quantum loop corrections induce the scale anomaly and generate the CW potential leading to the spontaneous scale-symmetry breaking. Since the CW potential merely includes a logarithmic field dependence, it exponentially suppresses the tunneling rate at low temperatures. This intrinsic feature naturally induces strong enough supercooling, hence is particularly relevant for the PBH formation Gouttenoire:2023naa ; Gouttenoire:2023pxh ; Baldes:2023rqv ; Conaci:2024tlc ; Arteaga:2024vde ; Salvio:2023ynn ; Salvio:2023blb ; Banerjee:2024cwv ; Kierkla:2025vwp .
@@
Very recently, the importance of what is called the “soft scale-breaking” mass term has also been emphasized in addressing the CW-type supercooling phase transition Zhang:2025kbu . This soft-scale breaking term can dramatically alter the prediction for PBH formation, due to violation of the exponential nucleation approximation. However, it still leaves the successful stochastic GW background predictions essentially as they stand in the case of the pure CW scenario Jaeckel:2016jlh ; Marzola:2017jzl ; Iso:2017uuu ; Ghorbani:2017lyk ; Baldes:2018emh ; Prokopec:2018tnq ; Brdar:2018num ; Marzo:2018nov ; DelleRose:2019pgi ; Ghoshal:2020vud ; Levi:2022bzt ; Kierkla:2022odc ; Salvio:2023qgb ; Zhang:2024vpp . No matter how small such a deformation is, the soft-scale breaking is still active even at low enough temperature – much lower than the critical temperature for the thermal phase transition ($T_{c}$), but as low as the critical temperature for the bubble nucleation ($T_{n}$). Thus this deformation possibility has provided a new testable link between PBH dark matter and GW signatures in the CW-type scenario Zhang:2025kbu .
@@
Here $M$ is a CW scalar field; the $m_{0}^{2}$ term corresponds to what is called the soft-scale breaking term; $v$ denotes the vacuum expectation value of $M$; $c_{1,2}$ are parameters to be determined at one-loop level given a model. The potential barrier around the origin can be generated when $m_{0}^{2}>0$ with $c_{1}>0$, much like the case of the CW thermal phase transition. As long as $m_{0}^{2}$ is small enough, this phase transition will be supercooled due to the approximate scale invariance around the origin of the potential.
@@
Actually, the potential form Eq.(1) is also the one that arises in low-energy chiral effective models of QCD and QCD-like theories, such as a Nambu-Jona-Lasinio (NJL) description in the mean-field approximation (MFA) with the bosonization of chiral quark bilinear: $M\sim\bar{q}q$. In there, the logarithmic term $\sim M^{4}\ln M$ is generated as a consequence of the quantum scale anomaly, which is also related to the renormalization group evolution of the dynamically induced quartic coupling for $M$. In medium with, e.g., finite baryon chemical potential $\mu_{B}$, this term would get corrections depending on $\mu_{B}$. The soft-scale breaking term $\sim m_{0}^{2}M^{2}$, at $T\sim T_{n}\ll T_{c}$, would be governed by the vacuum contribution, which induces $m_{0}^{2}<0$ as the dynamical chiral symmetry breaking, and also Debye screening mass in medium, $\sim\mu_{B}^{2}M^{2}$, which drives $m_{0}^{2}$ to be positive. Furthermore, beyond the SM contribution, if coupled to the QCD chiral phase transition, might also come into play as the soft-scale breaking term. Thus, the chiral phase transition in QCD-like theories can also be argued in terms of the same class of the scale violation, the soft-scale breaking $\sim M^{2}$ and the quantum scale anomaly $\sim M^{4}\ln M$, as discussed in Zhang:2025kbu .
@@
In QCD of the SM, which we shall call ordinary QCD, the chiral phase transition goes like crossover at around vanishing baryon chemical potential Aoki:2006we ; Borsanyi:2020fev ; Steinbrecher:2018phh . Thereby, a first-order and supercooling phase transition would be unlikely to take place within ordinary QCD framework. This feature is also reflected in the MFA-NJL description (see, e.g., reviews Hatsuda:1994pi ). The deviation from the CW-type phase transition is tightly linked with the fact that ordinary QCD generates a negatively too large soft-scale breaking mass ($m_{0}^{2}<0$ in Eq.(1)), as will be more explicitized later.
@@
In the literature Shao:2024dxt ; Shao:2024ygm , the possibility of realizing the supercooling from the QCD phase transition has been explored based on a couple of low-energy effective models. There it has been clarified that with a baryon chemical potential much higher than the typical QCD scale $(\lesssim 200-400\,{\rm MeV})$, a small phase transition rate $(\beta/H)$ enough to produce nano hertz GWs can be achieved. This, in terms of the soft-scale breaking, occurs due to not merely the creation of the Fermi surface, but the Debye mass $\sim\mu_{B}^{2}M^{2}$ that is positive and still large enough even at $T\sim T_{n}$.
@@
Another Debye mass effect as a positive soft-scale breaking has also been discussed in the recent literature Zhang:2025kbu . In the reference, in a context different from ordinary QCD, the chiral (axial) chemical potential $\mu_{5}$ in a model of scale-invariant dark QCD does the job along with mixing with an axial charged axionlike particle (ALP).
@@
In this paper, we first recap and clarify how the chiral phase transition in QCD-like theories can be CW-type with Beyond the SM. The key phenomenon necessary to be triggered is a cancellation of the soft-scale breaking terms at around $T=T_{c}$. We discuss an explicit dynamics to realize this key condition, which includes a well-known feature of many-flavor (walking) dark QCD and the recently addressed chiral phase transition with a large $\mu_{B}$ $\sim 1$ GeV in QCD Shao:2024dxt ; Shao:2024ygm , where for the latter, the Affeck-Dine baryogenesis Affleck:1984fy ; Linde:1985gh ; Dine:2003ax is assumed to leave a large amount of $\mu_{B}$ in the QCD phase transition epoch.
The CW-type supercooling phase transition can be realized even in QCD-like theories, where the phase transition essentially takes the potential form in Eq.(1). This is possible as long as the quantum scale anomaly takes the four- or three-dimensional form. The strongly magnetized medium is, therefore, still nontrivial because the four-dimensional system is to be reduced to the two-dimensional one characterized as $SO(2)\times SO(1,1)$ symmetry. This case will be disregarded in the present study.
We employ a two-flavor NJL model in the MFA, with monitoring the QCD case, and analyzing the thermal chiral phase transition in the presence of the chiral chemical potential $\mu_{5}$, nonzero strong CP phase $\theta$, in addition to the baryon chemical potential $\mu_{B}$. The scale violation classification is also operative in medium characterized by other chemical potentials, such as the isospin chemical potential ($\mu_{I}$), as long as the extreme condition is supplied via external gauging of the global chiral symmetry.
The three-flavor QCD case would have another driving force of the trilinear term $M^{3}$, which can also generate a potential barrier, associated with the $U(1)$ axial anomaly and the QCD instanton Kobayashi:1970ji ; Kobayashi:1971qz ; tHooft:1976snw . This cubic term, however, is not essential for the spontaneous breaking of the chiral symmetry – which does not destabilize the symmetric phase at $M=0$ and acts as higher order correction, when the negative $M^{2}$ term is present – and does not get thermal corrections from quark loop at the MFA level. Therefore, this cubic term would be subdominant when the CW-type chiral phase transition is realized and the supercooling is processed.
@@
  • In the NJL-like models, the several types of sources for “soft” scale violation, up to temperature $T$, are introduced as
    i) charge or matter density via gauging global currents (in part); e.g. $\mu_{B}$, $\mu_{5}$, the Polyakov loop (as an imaginary chemical potential),
@@
    iii) dynamical chiral symmetry breaking via four-fermion interactions.
  • At high enough $T$, the thermal correction cancels the CW-type scale anomaly of the form $\sim M^{4}\log M$, and the chiral phase transition is triggered simply by the generated Debye mass for $M$ and thus goes like the second-order type.
@@
  • If the scale symmetry of a theory is only logarithmically broken, like CW-type models, the high-$T$ expansion is insufficient to see the thermal phase transition, which is because the transition can take place at relatively lower $T$ in comparison with the case where a soft scale-breaking, negative mass-squared term $(\sim-M^{2})$ is present. In that case, the logarithmic term is not perfectly canceled enough to interfere creation of the potential barrier around $M=0$. Thus this type of theory undergoes a first-order phase transition, just like the CW-type models.
@@
  • In comparison with the CW-type model, QCD-like theories induce the dynamical chiral symmetry breaking, which reflects the negative mass-squared term for $M$ in the MFA-NJL description or bosonization. This deepens the curvature around $M=0$, so that the transition temperature becomes high enough that the high-$T$ expansion works well and cancels the net logarithmic term to wash out the barrier wall. Thus this causes the second-order, or crossover phase transition. This is why the first order is not realized along the $T$-axis in QCD-like theories. This also implies a key understanding of how the electroweak phase transition becomes crossover in the Standard Model with the negative soft-scale breaking term $\sim-H^{2}$.
@@
  • Going along the $\mu_{B}$ axis at $T=0$, the CW-type scale anomaly of the form $\sim M^{4}\log M$ is also canceled by the finite density ($\mu_{B}$) correction, in a way similar to the high $T$ correction case. Still, the finite $\mu_{B}$ generates not only the Debye mass for $M$, but also a potential barrier due to the creation of the Fermi surface which acts as a repulsive force. Therefore, at $T=0$, the chiral phase transition is of first order which is not generically of the CW type form. However, at finite $T$, it is possible to have the CW-type phase transition with an appropriate $\mu_{B}$ by controlling the net soft-scale breaking mass term.
[…]
@@
This paper is structured as follows: in Sec. II, we present the preliminary setup of the two-flavor NJL model and derive the thermodynamic potential in the MFA with nonzero $\mu_{5},\mu_{B},\theta$ at finite $T$. In Sec. III, we clarify the generic properties of the thermal chiral phase transition at $\mu_{B}=0$ in comparison with the CW-type phase transition, particularly paying our attention to the scale violation classes, without $\mu_{B}$ (Sec. III.1) and with $\mu_{B}$ (Sec. III.2). Then, in Sec. IV, we propose the ALP-assisted CW-type chiral supercooling and present the phenomenological constraints on the predicted ALP and discuss its experimental and cosmological detectability. The conclusion of this paper is provided in Sec. V.

### v4

@@ Abstract
We propose a new scenario to realize the Coleman-Weinberg (CW) type chiral phase transition in the QCD thermal history. This scenario predicts a heavy axionlike particle (ALP) with mass $\sim$ 5 MeV, consistently with the current experimental and cosmological bounds. The chiral phase transition is evaluated by monitoring ordinary QCD setup in a view of a two-flavor Nambu-Jona-Lasinio model including a simplified meson fluctuation contribution. The present work thus can open a new window to search for the ALP associated with the QCD phase transition epoch of the thermal history. The new QCD cosmological scenario potentially predicts rich epochs around the QCD scale: a mini-inflation; a nonperturbative preheating and/or reheating, which can provide characteristic gravitational wave and primordial black hole productions. This proposal is based on a generic classification of the order of the chiral phase transition at the level of the mean field approximation in view of the scale violation classes: the soft-scale breaking term and the CW-type scale anomaly term, in or off the medium with or without chemical potentials. On this theoretical ground, we also revisit existing scenarios which undergo the supercooling chiral phase transition, such as nearly scale-invariant QCD and QCD with a large baryon chemical potential.
@@ I Introduction
First order phase transitions or supercooling phase transitions currently involve broader interests over theoretical particle physics and cosmology Mazumdar:2018dfl . A particular motivation can be seen in cosmological applications to the thermal history of the Universe: the first-order/supercooling phase transition serves as productions for stochastic gravitational wave (GW) background and primordial black hole (PBH) formation Khlopov:2008qy ; Jung:2021mku ; Kodama:1982sf ; Hall:1989hr ; Kusenko:2020pcg ; Hashino:2021qoq ; Hashino:2022tcs ; He:2022amv ; Liu:2021svg ; Kawana:2022olo ; Lewicki:2023ioy ; Gouttenoire:2023naa ; Gouttenoire:2023pxh ; Salvio:2023blb ; Conaci:2024tlc ; Banerjee:2023brn ; Banerjee:2023vst ; Banerjee:2023qya ; Lewicki:2024ghw ; Flores:2024lng ; Cai:2024nln ; Florentino:2024kkf ; Goncalves:2024vkj ; Banerjee:2024fam ; Shao:2024dxt ; Banerjee:2024cwv ; Wu:2024lrp ; Hashino:2025fse ; Murai:2025hse ; Franciolini:2025ztf ; Kierkla:2025vwp ; Kanemura:2024pae as nonparticle dark matter candidate Belotsky:2018wph ; Carr:2020gox ; Baldes:2023rqv ; Balaji:2024rvo ; Jinno:2023vnr ; Zhang:2025kbu ; Kanemura:2024pae . It has also been argued that the supercooling electroweak or QCD phase transition can provide a possible initial condition for slow-roll inflation Zhang:2023acu ; Liu:2024xrh and subsequent particle production Wang:2022ygk ; Wang:2024tda .
@@
One typical modeling to access those cosmological applications would be a scale-symmetry phase transition via the Coleman-Weinberg (CW) mechanism Coleman:1973jx , assuming an appropriate presence of scalar field (presumably allowing coupled to the Standard Model (SM) plasma). All dimensionful parameters are absent at tree level, but quantum loop corrections induce the scale anomaly and generate the CW potential leading to spontaneous scale-symmetry breaking. Since the CW potential merely includes a logarithmic field dependence, it exponentially suppresses the tunneling rate at low temperatures. This intrinsic feature naturally induces strong enough supercooling, therefore, is particularly relevant for PBH formation Gouttenoire:2023naa ; Gouttenoire:2023pxh ; Baldes:2023rqv ; Conaci:2024tlc ; Arteaga:2024vde ; Salvio:2023ynn ; Salvio:2023blb ; Banerjee:2024cwv ; Kierkla:2025vwp .
@@
Very recently, the importance of what is called the “soft scale-breaking” mass term has also been emphasized in addressing the CW-type supercooling phase transition Zhang:2025kbu . This soft-scale breaking term can dramatically alter the prediction for PBH formation, due to violation of the exponential nucleation approximation. However, it still leaves the successful stochastic GW background predictions essentially as they are in the case of the pure CW scenario Jaeckel:2016jlh ; Marzola:2017jzl ; Iso:2017uuu ; Ghorbani:2017lyk ; Baldes:2018emh ; Prokopec:2018tnq ; Brdar:2018num ; Marzo:2018nov ; DelleRose:2019pgi ; Ghoshal:2020vud ; Levi:2022bzt ; Kierkla:2022odc ; Salvio:2023qgb ; Zhang:2024vpp . No matter how small such a deformation is, the soft-scale breaking is still active even at a low enough temperature – much lower than the critical temperature for the thermal phase transition ($T_{c}$), but as low as the critical temperature for the bubble nucleation ($T_{n}$). Thus this deformation possibility has provided a new testable link between PBH dark matter and GW signatures in the CW-type scenario Zhang:2025kbu .
@@
Here $M$ is a CW scalar field; the $m_{0}^{2}$ term corresponds to what is called the soft-scale breaking term; $v$ denotes the vacuum expectation value of $M$; $c_{1,2}$ are parameters to be determined at one-loop level given a model. The potential barrier around the origin can be generated when $m_{0}^{2}>0$ with $c_{1}>0$, much like in the case of the CW-type thermal phase transition. As long as $m_{0}^{2}$ is small enough, this phase transition will be supercooled due to the approximate scale invariance around the origin of the potential.
@@
Actually, the potential form Eq.(1) is also the one that arises in low-energy chiral effective models of QCD and QCD-like theories; e.g., a Nambu-Jona-Lasinio (NJL) description in the mean-field approximation (MFA) with the bosonization of chiral quark bilinear: $M\sim\bar{q}q$ and quark meson models with quark and meson fluctuation corrections included Mocsy:2004ab ; Khan:2011wai ; Zacchi:2017ahv . In there, the logarithmic term $\sim M^{4}\ln M$ is generated as a consequence of the quantum scale anomaly, which is also related to the renormalization group evolution of the (dynamically induced) quartic coupling for $M$. In medium with, e.g., finite baryon chemical potential $\mu_{B}$, this term would get corrections depending on $\mu_{B}$. The soft-scale breaking term $\sim m_{0}^{2}M^{2}$, at $T\sim T_{n}\ll T_{c}$, would be governed by the vacuum contribution, which induces $m_{0}^{2}<0$ as the spontaneous chiral symmetry breaking, and also Debye screening mass in the medium, $\sim\mu_{B}^{2}M^{2}$, which drives $m_{0}^{2}$ to be positive. Thus, the chiral phase transition in QCD-like theories can also be argued in terms of the same class of scale violation, soft-scale breaking $\sim M^{2}$, and the quantum scale anomaly $\sim M^{4}\ln M$, as discussed in Zhang:2025kbu .
@@
In QCD of the SM, which we shall call ordinary QCD, the chiral phase transition is a crossover around the vanishing baryon chemical potential Aoki:2006we ; Borsanyi:2020fev ; Steinbrecher:2018phh . Thereby, a first-order and supercooling phase transition would be unlikely to take place within the ordinary QCD framework. This feature is also reflected in the NJL-MF description (see, e.g., reviews Hatsuda:1994pi ) and quark-meson models (e.g., Mocsy:2004ab ; Khan:2011wai ; Zacchi:2017ahv for the case including both quark and meson fluctuations). As will be more explicitized later, the deviation from the CW-type phase transition is tightly linked with the fact that ordinary QCD generates a negatively too large soft-scale breaking mass ($m_{0}^{2}<0$ in Eq.(1)).
@@
In the literature Shao:2024dxt ; Shao:2024ygm , the possibility of realizing the supercooling from the QCD phase transition has been explored based on a couple of low-energy effective models. There, it has been clarified that with a large quark chemical potential of ${\cal O}(400\,{\rm MeV})$, a small phase transition rate $(\beta/H)$ can be realized and produce nano hertz GWs. This, in terms of soft-scale breaking, occurs due to not merely the creation of the Fermi surface, but the Debye mass $\sim\mu_{B}^{2}M^{2}$ that is positive and still large enough even at $T\sim T_{n}$.
@@
Another Debye mass effect as a positive soft-scale breaking term has also been discussed in the recent literature Zhang:2025kbu . In the reference, in a context different from ordinary QCD, the chiral (axial) chemical potential $\mu_{5}$ in a model of scale-invariant dark QCD does the job along with mixing with an axial charged axionlike particle (ALP).
@@
In this paper, we first recap and clarify how the chiral phase transition in QCD-like theories can be CW-type with Beyond the SM. The key phenomenon is the cancellation of the soft-scale breaking terms around $T=T_{c}$. We discuss an explicit dynamics to realize this key condition, which includes a well-known feature of many-flavor (walking) dark QCD and the recently addressed chiral phase transition with a large $\mu_{B}$ $\sim 1$ GeV in QCD Shao:2024dxt ; Shao:2024ygm , where for the latter, the Affleck-Dine baryogenesis Affleck:1984fy ; Linde:1985gh ; Dine:2003ax is assumed to leave a large amount of $\mu_{B}$ in the QCD phase transition epoch.
The CW-type supercooling phase transition can be realized even in QCD-like theories, where the phase transition essentially takes the potential form in Eq.(1). This is possible as long as the quantum scale anomaly takes a four- or three-dimensional form. The strongly magnetized medium is, therefore, still nontrivial because the four-dimensional system is to be reduced to the two-dimensional one characterized as $SO(2)\times SO(1,1)$ symmetry. This case will be disregarded in the present study.
We employ a two-flavor NJL model in the MFA, with monitoring the QCD case, and analyzing the thermal chiral phase transition in the presence of the chiral chemical potential $\mu_{5}$, nonzero strong CP phase $\theta$, in addition to the baryon chemical potential $\mu_{B}$. The scale violation classification is also operative in the medium characterized by other chemical potentials, such as the isospin chemical potential ($\mu_{I}$), as long as the extreme condition is supplied by external gauging of the global chiral symmetry.
The three-flavor QCD case would have another driving force of the trilinear term $M^{3}$, which can also generate a potential barrier, associated with the $U(1)$ axial anomaly and the QCD instanton Kobayashi:1970ji ; Kobayashi:1971qz ; tHooft:1976snw . This cubic term, however, is not essential for the spontaneous breaking of the chiral symmetry – which does not destabilize the symmetric phase at $M=0$ and acts as a higher order correction, when the negative $M^{2}$ term is present – and does not get thermal corrections at the quark-and/or meson one-loop level. Therefore, this cubic term would be subdominant when the CW-type chiral phase transition is realized and the supercooling is processed.
@@
  • In NJL-like models or quark meson-like models, the several types of sources for “soft”-scale violation, up to temperature $T$, are introduced as
    i) charge or matter density by gauging global currents (in part); e.g., $\mu_{B}$, $\mu_{5}$, the Polyakov loop (as an imaginary chemical potential),
@@
    iii) dynamical chiral symmetry breaking via four-fermion interactions and/or negative mass square for the chiral-order parameter-field $M$.
  • At high enough $T$, the thermal correction destructively cancels the CW-type scale anomaly of the form $\sim M^{4}\log M$ for the quark loop at vacuum, while for meson loops, the thermal correction is constructive, so the CW-type scale anomaly survives.
@@
  • If the scale symmetry at the vacuum is only logarithmically broken, like CW-type models, the high-$T$ expansion is insufficient to see the thermal phase transition, which is because the transition can take place at relatively lower $T$ in comparison with the case where a soft scale-breaking, negative mass-squared term $(\sim-M^{2})$ is present. In that case, even the fermion-loop induced CW-type scale anomaly is not perfectly canceled enough to interfere the creation of the potential barrier around $M=0$.
@@
  • Compared with the CW-type model, NJL-MF theories induce spontaneous chiral symmetry breaking, which reflects the negative mass-squared term for $M$. This deepens the curvature around $M=0$, so that the transition temperature becomes high enough that the high-$T$ expansion works well and cancels the net logarithmic term to wash out the barrier wall. Thus, this causes the second-order, or crossover phase transition.
@@
  • With nonzero $\mu_{B}$ at $T=0$, in NJL-MF models, the CW-type scale anomaly of the form $\sim M^{4}\log M$ is also canceled by the finite density ($\mu_{B}$) correction, in a way similar to the high $T$ correction case. Still, the finite $\mu_{B}$ generates not only the Debye mass for $M$, but also a potential barrier due to the creation of the Fermi surface which acts as a repulsive force. Therefore, at $T=0$, the chiral phase transition is of first order, which is not generically of the CW type form. However, at finite $T$, it is possible to have the CW-type phase transition with an appropriate $\mu_{B}$ by controlling the net soft-scale breaking mass term.
Passing through the clarification of the scale violation classes as above, we propose a new QCD cosmology scenario: the ALP-assisted CW-type chiral supercooling. The desired cancellation of the soft-scale breaking terms is achieved around $T=T_{c}$ by the portal coupling of the QCD-chiral order-parameter to an ALP field, which contributes significantly along with the chiral chemical potential, $\mu_{5}$, induced by the vacuum transition of the QCD sphaleron. It turns out that this scenario predicts rich and new QCD cosmology, which possibly includes a mini-inflation; preheating; reheating, and GW and/PBH production. As $T$ cools below as much as or less than $T_{n}$, where $\mu_{5}$ goes away due to the QCD sphaleron decoupling, in the chiral broken/confinement phase, we observe a massive ALP like a light axion, and the mass is constrained to be $\sim 5$ MeV, so as to survive the current experimental and cosmological bounds on the photon coupling $g_{a\gamma\gamma}$.
@@
This paper is structured as follows: in Sec. II, we present the preliminary setup of the two-flavor NJL model and derive the thermodynamic potential in the MFA with nonzero $\mu_{5},\mu_{B},\theta$ at finite $T$. In Sec. III, we clarify the generic properties of the thermal chiral phase transition at $\mu_{B}=0$ in comparison with the CW-type phase transition, particularly paying attention to the scale violation classes, without $\mu_{B}$ (Sec. III.1) and with $\mu_{B}$ (Sec. III.2). Then, in Sec. IV, we propose the ALP-assisted CW-type chiral supercooling and present the phenomenological constraints on the predicted ALP and discuss its experimental and cosmological constraints. The conclusion of this paper is provided in Sec. V.